http://www.jlab.org/~adaq/halog/html/0811_archive/081124151138.html
The problem was at the patch panel just before the ADC (the two cables had been flip-flopped). They are now plugged into the correct ADC channels.
This undoubtedly happened during testing, but I'm not sure when.
A copy of this log entry has been emailed to: meziani